### v2

# Gateformer: Advancing Multivariate Time Series Forecasting through Temporal and Variate-Wise Attention with Gated Representations

Abstract
There has been a recent surge of interest in time series modeling using the Transformer architecture. However, forecasting multivariate time series with Transformer presents a unique challenge as it requires modeling both temporal (cross-time) and variate (cross-variate) dependencies. While Transformer-based models have gained popularity for their flexibility in capturing both sequential and cross-variate relationships, it is unclear how to best integrate these two sources of information in the context of the Transformer architecture while optimizing for both performance and efficiency. We re-purpose the Transformer architecture to effectively model both cross-time and cross-variate dependencies. Our approach begins by embedding each variate independently into a variate-wise representation that captures its cross-time dynamics, and then models cross-variate dependencies through attention mechanisms on these learned embeddings. Gating operations in both cross-time and cross-variate modeling phases regulate information flow, allowing the model to focus on the most relevant features for accurate predictions. Our method achieves state-of-the-art performance across 13 real-world datasets and can be seamlessly integrated into other Transformer-based and LLM-based forecasters, delivering performance improvements up to 20.7% over original models. Code is available at this repository: https://github.com/nyuolab/Gateformer.
1 Introduction
Time series forecasting is a ubiquitous problem with essential applications in various domains, such as traffic [Lv et al., 2014], energy [Zhu et al., 2023], weather [Angryk et al., 2020], healthcare [Kaushik et al., 2020], and finance [Chen et al., 2012]. In challenging multivariate time series, accurate predictions rely on effectively modeling both intra-series (cross-time) and inter-series (cross-variate) dependencies [Zhang and Yan, 2023]. Transformer architecture [Vaswani et al., 2017] revolutionized the modern natural language processing due to their ability to efficiently process sequential data and capture semantic relationships over long sequence lengths. Naturally, there has been a huge interest in transfering the capabilities of the attention mechanism to the time series domain [Wen et al., 2023].
Most Transformer-based forecasters [Kitaev et al., 2020, Li et al., 2021, Liu et al., 2022b] naively embed data points from all variables at the same time step into point tokens and model intra-series dependencies using attention mechanisms. However, these individual points lack explicit semantic information, often leading to ineffective temporal dependency modeling. As a result, simpler linear models can outperform these Transformer-based methods [Zeng et al., 2023]. PatchTST [Nie et al., 2023] addresses this limitation by expanding the receptive field, dividing time series into sequences of patches, and applying cross-time attention among patches to capture temporal dependencies, yielding promising results. However, these methods still overlook critical cross-variate relationships. To explicitly model cross-variate correlations, methods such as iTransformer [Liu et al., 2024b] have been proposed. However, iTransformer embeds each series into a coarse variate representation through linear projection, overlooking intra-series dependencies.
In this work, we propose Gateformer, a Transformer-based model designed to explicitly capture both intra-series (cross-time) and inter-series (cross-variate) correlations. In the cross-time dependency modeling phase, we independently encode intra-series correlations of each variate into variate representations. These representations are then integrated with global temporal embeddings captured by an MLP through a gating mechanism. This gating operation regulates information flow, producing robust variate representations that serve as the foundation for subsequent cross-variate dependency modeling.
To capture cross-variate dependencies, we apply self-attention to these variate representations, generating variate-interacting embeddings. Recent work [Han et al., 2024] has shown that modeling cross-variate relationships enhances model capacity but requires larger datasets for effective correlation capture. Consequently, approaches focused on cross-variate modeling, like iTransformer, excel on high-dimensional datasets but show moderate performance on low-dimensional ones. To address this limitation, variate-interacting embeddings are integrated with outputs from the cross-time modeling phase (embeddings without variate interaction) through a gate operation to dynamically regulate the degree of cross-variate correlation, ensuring robust performance across datasets of varying sizes. Our main contributions in this work can be summarized as follows:
  • We propose Gateformer, which utilizes temporal and variate-wise attention with gated representations to enhance multivariate time series forecasting.
  • Our method achieves state-of-the-art results across 13 real-world benchmark datasets, ranking top-1 in 91 out of 130 settings and top-2 in 122 settings for varying prediction lengths and metrics.
  • The proposed framework integrates seamlessly with Transformer-based and LLM-based models, delivering up to 20.7% performance improvement over the original models. Its flexibility enables the development of foundational multivariate models based on this approach.
Figure 1: Transformer-based forecasters can be categorized based on their attention granularity (point-wise, patch-wise, and variate-wise) and their approach to modeling cross-time and cross-variate dependencies.
2 Related work
As shown in Figure 1, Transformer-based time series models can be categorized based on their granularity of representations (point-wise, patch-wise, and variate-wise) and their approach to modeling cross-time and cross-variate dependencies.
Point-wise Representations with Cross-time Attention
Many prior works embed multiple variates at the same timestamp into point tokens and apply attention mechanisms to capture temporal dependencies among them. Examples include Autoformer [Wu et al., 2021], FEDformer [Zhou et al., 2022], and Pyraformer [Liu et al., 2022b], which focus on optimizing the quadratic complexity of self-attention. Autoformer detects sub-series similarities with $O(L\log L)$ complexity, FEDformer leverages frequency-domain sparsity with $O(L)$ complexity, and Pyraformer uses pyramidal attention for short- and long-term dependencies with $O(L)$ complexity. However, these models often fail in multivariate forecasting due to the lack of explicit semantic meaning in point tokens and the loss of cross-variate correlations when merging all variables into single temporal tokens.
Patch-wise Representations with Cross-time Attention PatchTST [Nie et al., 2023] addresses the lack of local semantic information through patching. It divides time series into patches to expand the receptive field and applies self-attention to capture cross-time dependencies among patches. PatchTST processes each variate independently, allowing unique attention patterns to be learned for each series. This approach has consistently improved performance on benchmarks and inspired recent large-scale time series models [Zhou et al., 2023, Das et al., 2024, Jin et al., 2024, Liu et al., 2024c]. However, it focuses primarily on cross-time dependencies while neglecting critical cross-variate interactions.
Variate-wise Representations with Cross-variate Attention A notable example is iTransformer [Liu et al., 2024b], which expands the receptive field as an extreme form of patching. iTransformer encodes each variate’s time series into coarse variate-wise representations through linear projection and explicitly models cross-variate correlations among these representations. However, it does not account for cross-time (intra-series) dependencies.
Patch-wise Representations with Cross-time and Cross-variate Attention Crossformer [Zhang and Yan, 2023], a representative model, segments each variate’s series into a sequence of patches and captures both cross-time (intra-series) and cross-variate (inter-series) dependencies among them. However, modeling multivariate correlations at such a granular level can introduce unnecessary noise, potentially resulting in suboptimal forecasting performance.
Unlike previous works, our approach captures both cross-time and cross-variate correlations through variate-wise representations. To enhance the quality of these representations, we integrate gating mechanisms to regulate information flow, allowing the model to focus on the most relevant features for accurate predictions. More details are provided in Section 3.
Figure 2: Overall model architecture: The model encodes each variate’s series independently through two distinct pathways to obtain variate-wise representations: (1) temporal dependency embeddings that capture cross-time dependencies through patching and temporal-wise attention, and (2) global temporal embeddings that encode global temporal patterns through an MLP. These complementary embeddings are integrated through a gating operation to form variate embeddings, which serve as input for cross-variate dependency modeling. Variate-wise attention is then applied on variate embeddings to model multivariate dependencies, producing variate-interacting embeddings. Finally, a copy of the variate embeddings (without interaction) is combined with the variate-interacting embeddings through gating to regulate cross-variate correlations. The resulting output is then passed through a projection layer to generate predictions.
3 Methodology
3.1 Problem definition
Given multivariate historical observations $\mathbf{X}\in\mathbb{R}^{N\times T}$ , time series forecasting aims to predict the next $F$ time steps, denoted as $\mathbf{\hat{Y}}\in\mathbb{R}^{N\times F}$ , where $N$ is the number of variates (or channels) and $T$ is the look-back window length. The objective is to minimize the mean squared error between $\hat{\mathbf{Y}}$ and the ground truth $\mathbf{Y}\in\mathbb{R}^{N\times F}$.
3.2 Model Overview
As shown in Figure 2, the proposed approach leverages the Transformer framework to explicitly capture cross-time and cross-variate dependencies, which are essential for multivariate forecasting.
Cross-time Dependency Modeling Phase Each variate’s time series is processed independently through two distinct pathways to generate variate-wise representations: (1) Temporal Dependency Embeddings, which capture cross-time correlations using patching and temporal-wise attention; and (2) Global Temporal Embeddings, which encode global temporal patterns using an MLP. These complementary embeddings are integrated through a gating operation to form the final variate embeddings.
Cross-variate Dependency Modeling Phase Variate-wise attention is applied to the variate embeddings to capture multivariate correlations, resulting in variate-interacting embeddings. However, modeling cross-variate dependencies requires more training data and can struggle with smaller datasets. To address this limitation, we integrate the variate-interacting embeddings with variate-independent embeddings (a copy of outputs from the cross-time modeling phase) to regulate the degree of cross-variate correlation, which is especially crucial in the presence of data scarcity.
Output Projection Layer The output from the cross-variate modeling phase is passed through a linear layer to generate the final predictions.
3.3 Model Structure
3.3.1 Cross-time Dependency Modeling
Accurately capturing intra-series dependencies is essential for creating high-quality series representations, which serve as the basis for subsequent cross-series dependency modeling. To achieve this, we independently encode each variate’s series as a variate-wise (series-wise) token, motivated by the following:
  • Each series is processed independently through the self-attention module, enabling it to learn unique attention patterns.
  • This approach minimizes the risk of noise in one series affecting others.
Temporal Dependency Each series $\mathbf{x}^{(i)}\in\mathbb{R}^{T}$ is normalized using reversible instance normalization (RevIN) to achieve zero mean and unit standard deviation, mitigating distribution shifts between training and testing data [Kim et al., 2022]. To preserve local semantics and increase the receptive field, $\mathbf{x}^{(i)}$ is divided into non-overlapping patches $\mathbf{x}^{(i)}_{P}\in\mathbb{R}^{P\times L_{p}}$, where $P=\lfloor\frac{T-L_{p}}{S}\rfloor+2$ denotes the number of patches, $L_{p}$ is the patch size, and $S$ is the stride (distance between the starting points of consecutive patches). These patches are mapped to $\hat{\mathbf{x}}^{(i)}_{P}\in\mathbb{R}^{P\times d_{m}}$ through a trainable linear projector to match the Transformer’s hidden dimension $d_{m}$.
To model cross-time (intra-series) dependency, we perform self-attention over $\hat{\mathbf{x}}^{(i)}_{P}$ to model correlation between patches. Queries, keys, and values ($\mathbf{Q}^{(i)}_{P},\mathbf{K}^{(i)}_{P},\mathbf{V}^{(i)}_{P}\in\mathbb{R}^{P\times d_{k}}$) are computed through linear mappings on $\hat{\mathbf{x}}^{(i)}_{P}$, where $d_{k}$ is the projected dimension. The self-attention output $\mathbf{O}^{(i)}_{cross-time}\in\mathbb{R}^{P\times d_{k}}$ is obtained as:
$\mathbf{O}^{(i)}_{cross-time}=\textsc{Attention}(\mathbf{Q}^{(i)}_{P},\mathbf{K}^{(i)}_{P},\mathbf{V}^{(i)}_{P})$
$=\textsc{Softmax}\left(\frac{\mathbf{Q}^{(i)}_{P}\mathbf{K}^{(i)\top}_{P}}{\sqrt{d_{k}}}\right)\mathbf{V}^{(i)}_{P}$
At this stage, the temporal dependency is now captured in $\mathbf{O}^{(i)}_{cross-time}$. We then reshape $\mathbf{O}^{(i)}_{cross-time}$ to $\mathbf{O^{\prime}}^{(i)}_{cross-time}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}=P\times d_{k}$) and pass it through a feed-forward network to generate the temporal dependency embedding $\mathbf{v}^{(i)}_{T}\in\mathbb{R}^{d_{m}}$.
Global Temporal Pattern Although partitioning time series into patches and modeling dependencies among them effectively captures cross-time dependencies, as demonstrated by the success of PatchTST [Nie et al., 2023], it inevitably leads to information loss since patches only represent portions of the original series. To mitigate this, each series $\mathbf{x}^{(i)}$ is passed through a shared feed-forward network to extract global temporal patterns shared across variates and generate the global temporal embeddings. Specifically, the network linearly transforms $\mathbf{x}^{(i)}$ into $\mathbf{v}^{(i)}_{G}\in\mathbb{R}^{d_{m}}$. Recent studies [Zeng et al., 2023, Das et al., 2023, Li et al., 2023] have shown that MLPs can effectively capture temporal patterns often shared across different time series.
3.3.2 Gated Fusion
To improve the quality of variate embeddings, a gating mechanism is used to integrate representations encoding temporal dependencies $\mathbf{v}^{(i)}_{T}$ with those capturing global temporal patterns $\mathbf{v}^{(i)}_{G}$. This allows the model to focus on relevant information for accurate predictions. Formally, both embeddings are projected into a latent space, combined element-wise, and passed through a sigmoid function to compute the gate value. The embeddings are then integrated as follows:
$Gate=\sigma(\mathbf{v}^{(i)}_{T}\boldsymbol{W}_{g1}+\mathbf{v}^{(i)}_{G}\boldsymbol{W}_{g2})$ (1)
$\mathbf{s}^{(i)}=Gate\odot\mathbf{v}^{(i)}_{T}+(1-Gate)\odot\mathbf{v}^{(i)}_{G}$ (2)
where $\boldsymbol{W}_{g1}$ and $\boldsymbol{W}_{g2}$ are learnable parameters, $\sigma(\cdot)$ is the sigmoid function, and $\mathbf{s}^{(i)}\in\mathbb{R}^{d_{m}}$ is the final series representation. This mechanism produces a robust variate representation that encodes both temporal dependencies and global temporal patterns, providing a strong basis for cross-variate (cross-series) dependency modeling.
3.3.3 Cross-variates Dependency Modeling
A variate-wise attention module is used to model cross-variate dependencies. Given the variate embeddings $\mathbf{S}=\{\mathbf{s}^{(0)},\ldots,\mathbf{s}^{(N)}\}\in\mathbb{R}^{N\times d_{m}}$, the self-attention mechanism projects $\mathbf{S}$ into queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{m}}$. Following the standard self-attention procedure, the output $\mathbf{O}_{cross-variate}\in\mathbb{R}^{N\times d_{m}}$ is computed as:
$\mathbf{O}_{cross-variate}=\textsc{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})$
$=\textsc{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}$
While the multivariate correlations are encoded in $\mathbf{O}_{cross-variate}$, this approach requires more training data to effectively capture cross-variate dependencies, leading to moderate performance on smaller datasets. To mitigate this, a gating mechanism is used again to integrate the embeddings from channel-interacting $\mathbf{O}_{cross-variate}$ and channel-independent perspectives $\mathbf{S}$, ensuring robust forecasting results across both large and small datasets.
3.3.4 Output Projection Layer
Directly producing long-term forecasts using a linear layer has proven highly effective [Nie et al., 2023, Zeng et al., 2023]. Thus, a linear head is applied to $\mathbf{O}_{cross-variate}\in\mathbb{R}^{N\times d_{m}}$ to generate the prediction $\mathbf{\hat{Y}}\in\mathbb{R}^{N\times F}$. The mean and standard deviation used to normalize the input series are restored to the output predictions. Mean squared error (MSE) is used to measure the discrepancy between predictions $\mathbf{\hat{Y}}$ and ground truth $\mathbf{Y}$.
Table 1: Statistics of the thirteen datasets used in our experiments
Datasets Variates Timesteps Frequency
ETTh1&2 7 17,420 hourly
ETTm1&2 7 69,680 15 min
Traffic 862 17,544 hourly
Electricity 321 26,304 hourly
Weather 21 52,696 10 min
Exchange 8 7,588 daily
Solar-Energy 137 52,560 10 min
PEMS03 137 26,208 5 min
PEMS04 358 16,992 5 min
PEMS07 307 28,224 5 min
PEMS08 883 17,856 5 min
Table 2: Multivariate forecasting results with prediction lengths $T\in\{96,192,336,720\}$ and fixed look-back length $L=96$. Results are averaged from all prediction lengths. A lower value indicates better performance. Full results are listed in the Appendix G.
Models Gateformer iTransformer PatchTST Crossformer FEDformer Autoformer Stationary TimesNet SCINet DLinear
(Ours) [2024b] [2023] [2023] [2022] [2021] [2022c] [2023] [2022a] [2023]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETT(Avg) 0.373 0.394 0.383 0.399 0.381 0.397 0.685 0.578 0.408 0.428 0.465 0.459 0.471 0.464 0.391 0.404 0.689 0.597 0.442 0.444
Weather 0.253 0.276 0.258 0.278 0.259 0.281 0.259 0.315 0.309 0.360 0.338 0.382 0.288 0.314 0.259 0.287 0.292 0.363 0.265 0.317
Electricity 0.176 0.267 0.178 0.270 0.205 0.290 0.244 0.334 0.214 0.327 0.227 0.338 0.193 0.296 0.192 0.295 0.268 0.365 0.212 0.300
Traffic 0.412 0.276 0.428 0.282 0.481 0.304 0.550 0.304 0.610 0.376 0.628 0.379 0.624 0.340 0.620 0.336 0.804 0.509 0.625 0.383
Exchange 0.326 0.386 0.360 0.403 0.367 0.404 0.940 0.707 0.519 0.429 0.613 0.539 0.461 0.454 0.416 0.443 0.750 0.626 0.354 0.414
Solar-Energy 0.224 0.258 0.233 0.262 0.270 0.307 0.641 0.639 0.291 0.381 0.885 0.711 0.261 0.381 0.301 0.319 0.282 0.375 0.330 0.401
PEMS(Avg) 0.070 0.169 0.072 0.173 0.104 0.211 0.103 0.195 0.137 0.253 0.346 0.431 0.089 0.193 0.091 0.193 0.072 0.174 0.133 0.252
$1^{st}$ Count 7 7 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Table 3: Ablation study of model components across eight datasets, using prediction length $T=96$ and look-back window $L=96$.
Models ETTh1 ETTm1 ETTh2 ETTm2 Exchange Electricity Traffic Weather Avg. rank
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Gateformer 0.383 0.398 0.320 0.360 0.306 0.351 0.176 0.260 0.081 0.199 0.146 0.238 0.390 0.261 0.168 0.208 1.375 1.125
w/o Temporal-wise Attn. 0.393 0.404 0.327 0.361 0.325 0.360 0.179 0.261 0.084 0.203 0.147 0.239 0.395 0.266 0.176 0.217 3.25 3.000
w/o Gate in Variate-wise Attn. 0.385 0.400 0.327 0.362 0.334 0.367 0.182 0.264 0.082 0.200 0.144 0.239 0.387 0.261 0.168 0.208 2.250 2.375
w/o Global Temporal Embeddings 0.381 0.397 0.323 0.362 0.343 0.367 0.179 0.262 0.084 0.200 0.146 0.241 0.392 0.261 0.172 0.212 2.750 2.500
4 EXPERIMENTS
We extensively evaluate the proposed model on both long-term and short-term time series forecasting tasks, comparing its performance against nine state-of-the-art (SOTA) methods. An ablation study highlights the contribution of each model component. Additionally, our experiments demonstrate that the proposed method enhances both Transformer-based and LLM-based forecasters, delivering substantial performance improvements over the original models.
Dataset We use thirteen real-world datasets in the experiments, including Traffic, Electricity, Weather, Exchange, ETT (4 subsets) from Autoformer [Wu et al., 2021], Solar-Energy from LSTNet [Lai et al., 2018], and PEMS (4 subsets) from SCINet [Liu et al., 2022a]. For a fair comparison, we follow the standard protocol [Wu et al., 2023, Liu et al., 2024b], splitting the datasets into training, validation, and test sets using ratios of 6:2:2 for ETT and PEMS datasets, and 7:1:2 for the others. The splits are strictly chronological to ensure no data leakage occurs. The characteristics of these datasets are shown in Table 1, with additional details provided in Appendix A.1.
Table 4: Performance improvements achieved by our framework. The results are averaged across four prediction lengths $T\in\{96,192,336,720\}$ with a look-back window of $L=96$. Relative reductions in MSE and MAE (Promotion) are reported. Full results are provided in the Appendix D.
Models Autoformer Flowformer GPT4TS
MSE MAE MSE MAE MSE MAE
ECL Original 0.227 0.338 0.267 0.359 0.217 0.308
+ Our framework 0.198 0.283 0.171 0.263 0.177 0.267
Promotion 12.8% 16.3% 36.0% 26.7% 18.4% 13.3%
Weather Original 0.338 0.382 0.286 0.308 0.279 0.297
+ Our framework 0.257 0.281 0.255 0.279 0.261 0.282
Promotion 24.0% 26.4% 10.8% 9.4% 6.4% 5.0%
Table 5: Transfer learning results. Gateformer is pre-trained on PEMS07 dataset and then transferred to other datasets. Best results are marked in bold, with prediction lengths $T\in\{3,6,12,24\}$ and look-back window $L=96$.
Models Gateformer Models Trained from Scratch on Target Dataset
Zero-shot Fine-tuning Gateformer iTransformer PatchTST Crossformer FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03 3 0.049 0.148 0.044 0.140 0.045 0.142 0.047 0.147 0.055 0.166 0.051 0.150 0.109 0.231
6 0.064 0.167 0.052 0.152 0.053 0.153 0.057 0.159 0.069 0.184 0.060 0.161 0.112 0.235
12 0.092 0.198 0.065 0.168 0.066 0.170 0.071 0.174 0.099 0.216 0.090 0.203 0.126 0.251
24 0.146 0.251 0.088 0.197 0.092 0.201 0.093 0.201 0.142 0.259 0.121 0.240 0.149 0.275
Avg 0.088 0.191 0.062 0.164 0.064 0.166 0.067 0.170 0.091 0.206 0.081 0.189 0.124 0.248
PEMS04 3 0.064 0.164 0.060 0.156 0.061 0.159 0.064 0.164 0.071 0.184 0.062 0.162 0.122 0.250
6 0.077 0.181 0.067 0.166 0.068 0.168 0.073 0.175 0.081 0.191 0.069 0.173 0.119 0.245
12 0.101 0.210 0.081 0.183 0.083 0.186 0.078 0.183 0.105 0.224 0.098 0.218 0.138 0.262
24 0.162 0.269 0.110 0.216 0.114 0.221 0.095 0.205 0.153 0.275 0.131 0.256 0.177 0.293
Avg 0.101 0.206 0.080 0.180 0.081 0.184 0.078 0.182 0.103 0.219 0.090 0.202 0.139 0.263
PEMS08 3 0.057 0.155 0.051 0.144 0.052 0.147 0.055 0.153 0.064 0.175 0.117 0.161 0.153 0.255
6 0.070 0.172 0.059 0.154 0.060 0.156 0.064 0.165 0.076 0.190 0.129 0.173 0.157 0.257
12 0.094 0.200 0.073 0.170 0.075 0.172 0.079 0.182 0.168 0.232 0.165 0.214 0.173 0.273
24 0.157 0.258 0.104 0.200 0.108 0.204 0.115 0.219 0.224 0.281 0.215 0.260 0.210 0.301
Avg 0.094 0.196 0.072 0.167 0.074 0.170 0.078 0.180 0.133 0.220 0.157 0.202 0.173 0.272
4.1 FORECASTING RESULTS
Baselines
In this section, we conduct extensive experiments to evaluate the forecasting performance of our proposed model together with advanced forecasters. We carefully choose nine well-acknowledged forecasting models as our benchmark, including (1) Transformer-based methods: Autoformer [Wu et al., 2021], FEDformer [Zhou et al., 2022], Stationary [Liu et al., 2022c], Crossformer [Zhang and Yan, 2023], PatchTST [Nie et al., 2023], iTransformer [Liu et al., 2024b]; (2) Linear-based methods: DLinear [Zeng et al., 2023]; and (3) TCN-based methods: SCINet [Liu et al., 2022a], TimesNet [Wu et al., 2023].
Experiment settings All models were evaluated under a consistent experimental setup, with a look-back window of $L=96$ and a fixed training duration of 10 epochs. The prediction lengths were $T\in\{3,6,12,24\}$ for the PEMS dataset; $T\in\{96,192,336,720\}$ for other datasets. Baseline results, except for the PEMS dataset, are from Liu et al. [2024b]. For the PEMS dataset, we specifically cite the results for prediction lengths $T\in\{12,24\}$ from Liu et al. [2024b]. All experiments were implemented in PyTorch [Paszke et al., 2019] and conducted on a single NVIDIA A100 40GB GPU. We reproduced the baseline models for comparison using the implementation from the TimesNet [Wu et al., 2023] repository.
Main results Table 2 presents multivariate forecasting results averaged across four prediction lengths, with best results highlighted in red and second-best in underlined. Lower MSE/MAE values indicate more accurate predictions. Our model consistently achieves state-of-the-art performance across all datasets, demonstrating its effectiveness in multivariate forecasting. Full results are provided in Appendix G. The comparison between Gateformer, iTransformer [Liu et al., 2024b], and Crossformer [Zhang and Yan, 2023] is noteworthy. Although iTransformer and Crossformer explicitly capture cross-variate dependencies, they underperform compared to our method. iTransformer focuses solely on capturing cross-variate correlations while neglecting cross-time dependency modeling, highlighting its limitation in encoding a variate’s time series through linear projection alone. Crossformer, similar to our method, divides each variable’s series into patches to model cross-time dependencies, followed by capturing cross-variate interactions across patches. However, its moderate performance suggests that modeling multivariate correlations at such a granular level may introduce unnecessary noise. In contrast, our method first encodes input series into variate-wise representations that integrate both cross-time dependencies and global temporal patterns. These representations are then used to model cross-variate dependencies. However, the flexibility to learn cross-variate relationships can be a double-edged sword, as it requires more training data to do so effectively [Nie et al., 2023, Han et al., 2024], leading to moderate performance on smaller datasets. While iTransformer and Crossformer struggle with low-dimensional datasets like ETT, our model performs on par with PatchTST [Nie et al., 2023], a representative channel-independent model that does not capture cross-variate dependencies. This strength can be attributed to our gating mechanism, which dynamically regulates the degree of cross-variate correlation, a capability that is especially critical in limited data scenarios. In summary, our approach excels at modeling both temporal and multivariate dependencies, making it highly effective for real-world time series forecasting tasks.
4.2 MODEL ANALYSIS
Ablation Study of Model Components To investigate the impact of each component in our method, we conducted ablation experiments on eight datasets: ETT (4 subsets), Weather, Traffic, Electricity, and Exchange. Specifically, we compared our model with three modified versions:
  • w/o Temporal-wise Attn. Removes the embeddings that encode temporal dependencies from the variate representation, leaving only embeddings that capture global temporal patterns.
  • w/o Gate in Variate-wise Attn. Removes the gating mechanism in cross-variate dependency modeling.
  • w/o Global Temporal Embeddings. Removes the embeddings that encode global temporal patterns, retaining only the embeddings that capture temporal dependencies in the variate representation.
As shown in Table 3, Gateformer (first row), utilizing both temporal-wise and channel-wise attention with gated representations, achieves the best overall performance. Notably, removing temporal-wise correlation (second row) leads to the poorest performance, underscoring its critical role. The primary performance advantage of our method over iTransformer stems from its ability to model temporal-wise dependencies. While removing the gating mechanism for cross-variate correlation (third row) improves performance slightly on high-dimensional datasets like Traffic and Electricity, where abundant training data allows effective learning of cross-dimensional dependencies, it significantly degrades performance on smaller datasets like ETT. Overall, our proposed method consistently delivers optimal results across both high-dimensional and low-dimensional datasets.
4.3 Framework Generalizability
Capturing cross-variate correlations is a key feature of our model. However, the quadratic complexity of self-attention becomes a computational bottleneck as the number of variates increases, a common challenge in real-world applications. Several prior works [Li et al., 2021, Wu et al., 2021, 2022, Zhou et al., 2022] have introduced advanced attention mechanisms to mitigate this issue. Since our model retains the standard Transformer architecture, it can seamlessly integrate these advanced attention modules from Transformer-based forecasters by substituting the conventional attention mechanism. We demonstrate this versatility by incorporating advanced attention mechanisms from models like Autoformer [Wu et al., 2021] and Flowformer [Wu et al., 2022], as well as extending our framework to LLM-based models such as GPT4TS [Zhou et al., 2023] to enhance their multivariate correlation modeling.
Experiments on the Weather and Electricity datasets (look-back window $L=96$, prediction lengths $T\in{96,192,336,720}$) show significant improvements over the original models: 19.9% on Autoformer, 20.7% on Flowformer, and 10.8% on GPT4TS, as shown in Table 4. These results demonstrate our framework’s effectiveness and adaptability across different model architectures. Detailed results are provided in Appendix D.
4.4 REPRESENTATION LEARNING
We evaluate the representations learned by our model, focusing on their generalization and transferability across datasets. The PEMS datasets are used as a benchmark, as they comprise four high-dimensional time series subsets, each containing a substantial amount of data to ensure robust and reliable evaluation results.
Transfer learning To assess representation transferability, we pre-trained the model on the largest dataset (PEMS07) for 10 epochs, then transferring it to PEMS03, PEMS04, and PEMS08 datasets. As shown in Table 5, the model’s zero-shot forecasting performance is comparable to most competitive baseline models. A key strength of our method lies in its ability to effectively capture cross-variate correlations, enabling superior generalization and transferability to unseen datasets. Furthermore, pre-training followed by fine-tuning for a few epochs consistently outperforms training solely on downstream datasets, demonstrating the value of transferable representations learned during pre-training.
Figure 3: Investigation of the memory-efficient training strategy: The memory footprint (right) is significantly reduced when training on randomly selected variates, while maintaining consistent results (left) across different sampling ratios.
4.5 MODEL EFFICIENCY
Memory-Efficient Training Strategy Our model captures multivariate correlations by applying the attention mechanism along the variate dimension. However, the quadratic complexity of the attention mechanism becomes a computational bottleneck when training models on high-dimensional datasets such as Traffic. To mitigate this, we implement a computationally efficient training strategy that randomly samples a subset of variates for each batch, allowing the model to train on only the selected variates. During inference, the model generates forecasts for all variates. This random sampling of variates acts as a form of regularization, enabling the model to learn robust and generalizable patterns. As shown in Figure 3, the forecasting performance remains stable across various sampling ratios, while the memory footprint is significantly reduced.
Model Efficiency To evaluate the efficiency of our model, we compare the forecasting accuracy, training time, and memory footprint of the following models: Gateformer, Gateformer with the memory-efficient training strategy, iTransformer [Liu et al., 2024b], PatchTST [Nie et al., 2023], Crossformer [Zhang and Yan, 2023], TimesNet [Wu et al., 2023], and DLinear [Zeng et al., 2023]. For a fair comparison, all models use the same hidden dimension and batch size on a representative high-dimensional dataset (Traffic).
As shown in Figure 4, in terms of training speed, the linear model (DLinear) is the fastest. Gateformer, which captures both cross-time and cross-variate dependencies, is slower than iTransformer, which models only multivariate correlations, and PatchTST, which focuses purely on cross-time correlations. However, by adopting an efficient training strategy (training on a randomly sampled 20% of variates in each batch while forecasting all variates), Gateformer (Efficient) achieves comparable training speed to iTransformer with a lower memory footprint and superior forecasting performance.
Figure 4: Model efficiency comparison on Traffic dataset.
5 CONCLUSION AND FUTURE WORK
We propose a method that employs temporal and variate-wise attention modules to effectively capture both cross-time and cross-variate correlations. Gating mechanisms are incorporated to regulate the flow of information, enabling the model to focus on the most relevant features for accurate predictions. Experimentally, our approach achieves state-of-the-art performance and demonstrates flexibility, delivering up to a 20.7% performance improvement when integrated with other Transformer-based forecasters. In the future, we aim to explore large-scale pre-training and extend our method to a wider range of time series analysis tasks.
Software and Data
All implementation details, including dataset descriptions, metrics, models, and experiment configurations, are provided in Appendix A. The code is submitted as Supplementary Material to the OpenReview system.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Experimental Details
A.1 Datasets
We assess long-term forecasting performance on ten widely used benchmarks, including the four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Electricity, Traffic, Exchange (from Wu et al. [2021]), and Solar-Energy (from Lai et al. [2018]). Additionally, we evaluate short-term forecasting on the four PEMS datasets (PEMS03, PEMS04, PEMS07, PEMS08) as used in Liu et al. [2022a]. A brief description of each dataset is provided below.
  • ETT Dataset comprises two sub-datasets: ETTh and ETTm, which were collected from electricity transformers. Data were recorded at 15-minute and 1-hour intervals for ETTm and ETTh, respectively, spanning from July 2016 to July 2018.
  • Weather Dataset includes 21 meteorological indicators recorded every 10 minutes during 2020.
  • Electricity Dataset contains hourly electricity consumption data for 321 customers from 2012 to 2014.
  • Traffic Dataset provides hourly road occupancy rates measured by various sensors on San Francisco Bay Area freeways, collected by the California Department of Transportation.
  • Exchange Dataset collects daily exchange rates from 8 countries, spanning from 1990 to 2016.
  • Solar-Energy records the solar power production of 137 PV plants in 2006, which are sampled every 10 minutes.
  • PEMS Dataset contains public traffic network data from California, recorded at 5-minute intervals.
We adopt the same data processing and train-validation-test split protocol as TimesNet [Wu et al., 2023], ensuring datasets are strictly divided chronologically to prevent data leakage. For forecasting, the lookback series length is fixed at $L=96$ across all datasets. Prediction lengths are set as follows: $T\in\{96,192,336,720\}$ for ETT, Weather, ECL, Traffic, Solar-Energy, and Exchange; $T\in\{3,6,12,24\}$ for PEMS. Detailed dataset information is provided in Table 6.
Table 6: Detailed dataset descriptions. The dimension refers to the number of variates in each dataset and the dataset size is organized in (training, validation, testing).
Dataset Dim. Prediction Length Dataset Size Frequency Information
ETTh1, ETTh2 7
{96, 192, 336, 720}
(8545, 2881, 2881) Hourly Electricity
ETTm1, ETTm2 7
{96, 192, 336, 720}
(34465, 11521, 11521) 15min Electricity
Exchange 8
{96, 192, 336, 720}
(5120, 665, 1422) Daily Economy
Weather 21
{96, 192, 336, 720}
(36792, 5271, 10540) 10min Weather
ECL 321
{96, 192, 336, 720}
(18317, 2633, 5261) Hourly Electricity
Traffic 862
{96, 192, 336, 720}
(12185, 1757, 3509) Hourly Transportation
Solar-Energy 137
{96, 192, 336, 720}
(36601, 5161, 10417) 10min Energy
PEMS03 358
{3, 6, 12, 24}
(15617, 5135, 5135) 5min Transportation
PEMS04 307
{3, 6, 12, 24}
(10172, 3375, 3375) 5min Transportation
PEMS07 883
{3, 6, 12, 24}
(16911, 5622, 5622) 5min Transportation
PEMS08 170
{3, 6, 12, 24}
(10690, 3548, 3548) 5min Transportation
A.2 Implementation Details
All experiments were implemented in PyTorch [Paszke et al., 2019] and conducted on a single NVIDIA A100 40GB GPU. We used MSE as the loss function, a batch size of 8, and the Adam optimizer. The initial learning rate was selected from $\{10^{-4},5\times 10^{-4},10^{-3}\}$, with training running for a maximum of 10 epochs and early stopping applied if validation loss did not improve within three epochs. The number of encoder blocks was chosen from $L\in\{1,2,3,4\}$, and the hidden state dimension was selected from $\{64,128,256,512,1024\}$. Evaluation metrics included mean square error (MSE) and mean absolute error (MAE), with results averaged over three random seeds. Baseline models were reproduced using the TimesNet [Wu et al., 2023] repository, following configurations from their respective original papers or official code.
Figure 5: Forecasting performance (MSE) across different look-back windows $L\in\{24,48,96,192,336,720\}$ on Traffic, Electricity, and Weather datasets, with prediction length $T=96$.
Appendix B Varying Look-back Window
In principle, expanding the look-back window increases the receptive field, which should enhance forecasting accuracy. However, previous studies [Zeng et al., 2023] have shown that most Transformer-based models fail to demonstrate this expected improvement, revealing their limitations in processing extended temporal sequences. Our model, in contrast, effectively utilizes the increased receptive field, achieving lower MSE scores with longer look-back windows, as demonstrated in Figure 5.
Appendix C Hyperparameter Sensitivity
To evaluate the sensitivity of our method to hyperparameter settings, we conducted experiments by varying the number of Transformer blocks $L\in\{1,2,3,4\}$, model dimensions $D\in\{128,256,512,1024\}$, and learning rates $lr\in\{10^{-4},3\times 10^{-4},5\times 10^{-4},10^{-3}\}$ on the ETTm1, Weather, Electricity, and Traffic datasets. As shown in Figure 6, the learning rate has the strongest impact on model performance and requires careful tuning. Model performance generally improves with larger hidden dimensions, while remaining relatively stable across different numbers of Transformer blocks.
Appendix D Framework Generalizability
As discussed in Section 4.3, our proposed framework can be seamlessly integrated into other Transformer-based and LLM-based models to consistently improve their performance, with full forecasting results shown in Table 7. Since our model explicitly captures cross-variate dependencies, the variate-wise attention becomes a computational bottleneck when training on high-dimensional datasets such as Electricity, Traffic, and Solar-Energy. To mitigate this, we replace the conventional quadratic-time attention mechanism with the linear-time attention proposed in Flowformer [Wu et al., 2022] and compare its performance against competitive baselines. These results are provided in Figure 7.
Table 7: Full results of performance improvements achieved using our proposed framework.
Models
Autoformer
Flowformer
GPT4TS
[2021]
[2022]
[2023]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
Electricity Original 96 0.201 0.317 0.215 0.320 0.197 0.290
192 0.222 0.334 0.259 0.355 0.201 0.292
336 0.231 0.338 0.296 0.383 0.217 0.309
720 0.254 0.361 0.296 0.380 0.253 0.339
Avg
0.227
0.338
0.267
0.359
0.217
0.308
+ Our framework 96 0.173 0.259 0.144 0.237 0.150 0.241
192 0.180 0.267 0.160 0.251 0.163 0.254
336 0.198 0.285 0.175 0.269 0.180 0.272
720 0.241 0.320 0.205 0.296 0.214 0.301
Avg
0.198
0.283
0.187
0.274
0.177
0.267
Weather Original 96 0.266 0.336 0.182 0.233 0.203 0.244
192 0.307 0.367 0.250 0.288 0.247 0.277
336 0.359 0.395 0.309 0.329 0.297 0.311
720 0.419 0.428 0.404 0.385 0.368 0.356
Avg
0.338
0.382
0.286
0.308
0.279
0.297
+ Our framework 96 0.175 0.218 0.171 0.213 0.175 0.214
192 0.219 0.258 0.220 0.258 0.228 0.260
336 0.281 0.300 0.275 0.297 0.282 0.301
720 0.353 0.350 0.353 0.349 0.359 0.352
Avg
0.257
0.281
0.255
0.279
0.261
0.282
Figure 6: Forecasting results (MSE) varying with the number of Transformer blocks, the model’s hidden dimension, and the learning rate. The results were collected with a prediction horizon of $T=96$ and a look-back window of $L=96$.
Figure 7: Performance comparison on high-dimensional datasets. Gateformer (Efficient) replaces the conventional quadratic-time attention mechanism with Flowformer’s optimized linear-time attention to improve computational efficiency. Prediction horizon $T=96$; look-back window $L=96$.
Appendix E Unified Cross-Dataset Training
Figure 8: Analysis of unified cross-dataset training: Models are pre-trained on four combinations of PEMS datasets with increasing training set sizes, then fine-tuned on target datasets. Prediction length $T=24$; look-back window $L=96$.
We selected the PEMS dataset as our benchmark as it contains four high-dimensional time-series subsets (PEMS03, PEMS04, PEMS07, and PEMS08), each with substantial data that ensures robust evaluation. Our model includes temporal-wise and variate-wise self-attention modules. The temporal-wise attention module captures intra-series dependencies for each series independently. The variate-wise attention module operates on the variate dimension, enabling unified training across datasets with varying numbers of variates. To investigate the benefits of data scaling, we trained our model on four progressively larger combined datasets: (1) PEMS08, (2) PEMS08 + PEMS04, (3) PEMS08 + PEMS04 + PEMS03, and (4) PEMS08 + PEMS04 + PEMS03 + PEMS07. These datasets were combined in ascending order of dimensionality. As shown in Figure 8, forecasting accuracy improved consistently across all PEMS datasets with larger training set sizes. This highlights our model’s ability to learn generalizable and transferable representations through unified training while preventing catastrophic forgetting, a common challenge in combined training Liu et al. [2024a]. Our results highlight the potential of building large-scale multivariate forecasting models trained across datasets using our framework.
Figure 9: Visualization of 96-step forecasting on Traffic dataset with look-back window $L=96$.
Figure 10: Visualization of 96-step forecasting on Electricity dataset with look-back window $L=96$. denotes models integrated with our framework.
Appendix F Visualization
We visualize long-term forecasting results of Gateformer against baseline models on the Traffic dataset in Figure 9, where Gateformer demonstrates superior prediction accuracy. Figure 10 compares the forecasting performance of Transformer-based models with and without integration of our framework on the Electricity dataset. Models integrated with our approach consistently show improved prediction accuracy.
Appendix G Full Multivariate Forecasting Results
We extensively evaluated our model’s performance against competitive baselines on well-acknowledged forecasting benchmarks. Table 8 presents comprehensive short-term forecasting results on the four PEMS subsets. Table 9 details long-term forecasting results across nine challenging benchmarks.
Table 8: Full short-term multivariate forecasting results with prediction lengths $T\in\{3,6,12,24\}$ and input length $L=96$. The best results are highlighted in red and the second bests are marked in blue.
Model
Gateformer
iTransformer
PatchTST
Crossformer
FEDformer
Autoformer
Stationary
TimesNet
SCINet
DLinear
[2024b]
[2023]
[2023]
[2022]
[2021]
[2022c]
[2023]
[2022a]
[2023]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
PEMS03
3
0.045
0.142
0.047 0.147
0.055
0.166
0.051
0.150
0.109
0.231
0.285
0.394
0.063
0.168
0.068
0.173
0.048
0.147
0.069
0.183
6
0.053
0.153
0.057
0.159
0.069
0.184
0.060
0.161
0.112
0.235
0.254
0.363
0.070
0.176
0.074
0.179
0.056 0.158
0.086
0.204
12
0.066
0.170
0.071
0.174
0.099
0.216
0.090
0.203
0.126
0.251
0.272
0.385
0.081
0.188
0.085
0.192
0.066
0.172
0.122
0.243
24
0.092 0.201
0.093
0.201
0.142
0.259
0.121
0.240
0.149
0.275
0.334
0.440
0.105
0.214
0.118
0.223
0.085
0.198
0.201
0.317
Avg
0.064
0.166
0.067
0.170
0.091
0.206
0.081
0.189
0.124
0.248
0.286
0.396
0.080
0.187
0.086
0.192
0.064
0.169
0.120
0.237
PEMS04
3
0.061
0.159
0.064
0.164
0.071
0.184
0.062
0.162
0.122
0.250
0.305
0.414
0.076
0.181
0.075
0.179
0.060
0.159
0.096
0.218
6
0.068
0.168
0.073
0.175
0.081
0.191
0.069
0.173
0.119
0.245
0.361
0.449
0.080
0.187
0.079
0.183
0.067
0.169
0.113
0.236
12
0.083
0.186
0.078 0.183
0.105
0.224
0.098
0.218
0.138
0.262
0.424
0.491
0.088
0.196
0.087
0.195
0.073
0.177
0.148
0.272
24
0.114
0.221
0.095 0.205
0.153
0.275
0.131
0.256
0.177
0.293
0.459
0.509
0.104
0.216
0.103
0.215
0.084
0.193
0.224
0.340
Avg
0.081
0.184
0.078 0.182
0.103
0.219
0.090
0.202
0.139
0.263
0.387
0.466
0.087
0.195
0.086
0.193
0.071
0.175
0.145
0.267
PEMS07
3
0.042
0.130
0.046
0.139
0.052
0.158
0.050
0.141
0.102
0.218
0.201
0.326
0.069
0.169
0.068
0.165
0.043 0.132
0.061
0.172
6
0.049
0.141
0.054
0.150
0.061
0.169
0.057
0.152
0.104
0.219
0.253
0.373
0.074
0.175
0.073
0.171
0.050 0.144
0.078
0.196
12
0.061
0.156
0.067 0.165
0.095
0.207
0.094
0.200
0.109
0.225
0.199
0.336
0.083
0.185
0.082
0.181
0.068
0.171
0.115
0.242
24
0.086
0.187
0.088 0.190
0.150
0.262
0.139
0.247
0.125
0.244
0.323
0.420
0.102
0.207
0.101
0.204
0.119
0.225
0.210
0.329
Avg
0.060
0.154
0.064 0.161
0.090
0.199
0.085
0.185
0.110
0.227
0.244
0.364
0.082
0.184
0.081
0.180
0.070
0.168
0.116
0.235
PEMS08
3
0.052
0.147
0.055 0.153
0.064
0.175
0.117
0.161
0.153
0.255
0.434
0.463
0.084
0.183
0.088
0.186
0.059
0.158
0.094
0.215
6
0.060
0.156
0.064 0.165
0.076
0.190
0.129
0.173
0.157
0.257
0.526
0.541
0.092
0.191
0.096
0.195
0.065
0.167
0.112
0.234
12
0.075
0.172
0.079 0.182
0.168
0.232
0.165
0.214
0.173
0.273
0.436
0.485
0.109
0.207
0.112
0.212
0.087
0.184
0.154
0.276
24
0.108
0.204
0.115 0.219
0.224
0.281
0.215
0.260
0.210
0.301
0.467
0.502
0.140
0.236
0.141
0.238
0.122
0.221
0.248
0.353
Avg
0.074
0.170
0.078 0.180
0.133
0.220
0.157
0.202
0.173
0.272
0.466
0.498
0.106
0.204
0.109
0.208
0.083
0.183
0.152
0.270
$1^{st}$ Count
14
16
0
0
0
0
0
0
0
0
0
0
0
0
0
0
8 5
0
0
Table 9: Full results of multivariate long-term forecasting with a fixed input length $L=96$ for all datasets. Baseline results are from Liu et al. [2024b]. The best results are highlighted in red and the second bests are marked in blue.
Model
Gateformer
iTransformer
PatchTST
Crossformer
FEDformer
Autoformer
Stationary
TimesNet
SCINet
DLinear
[2024b]
[2023]
[2023]
[2022]
[2021]
[2022c]
[2023]
[2022a]
[2023]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Traffic
96
0.390
0.261
0.395 0.268
0.462
0.295
0.522
0.290
0.587
0.366
0.613
0.388
0.612
0.338
0.593
0.321
0.788
0.499
0.650
0.396
192
0.409
0.270
0.417 0.276
0.466
0.296
0.530
0.293
0.604
0.373
0.616
0.382
0.613
0.340
0.617
0.336
0.789
0.505
0.598
0.370
336
0.424
0.278
0.433 0.283
0.482
0.304
0.558
0.305
0.621
0.383
0.622
0.337
0.618
0.328
0.629
0.336
0.797
0.508
0.605
0.373
720
0.459
0.296
0.467 0.302
0.514
0.322
0.589
0.328
0.626
0.382
0.660
0.408
0.653
0.355
0.640
0.350
0.841
0.523
0.645
0.394
Avg
0.421
0.276
0.428 0.282
0.481
0.304
0.550
0.304
0.610
0.376
0.628
0.379
0.624
0.340
0.620
0.336
0.804
0.509
0.625
0.383
Solar-Energy
96
0.182
0.222
0.203 0.237
0.234
0.286
0.310
0.331
0.242
0.342
0.884
0.711
0.215
0.249
0.250
0.292
0.237
0.344
0.290
0.378
192
0.227
0.257
0.233 0.261
0.267
0.310
0.734
0.725
0.285
0.380
0.834
0.692
0.254
0.272
0.296
0.318
0.280
0.380
0.320
0.398
336
0.242
0.274 0.248
0.273
0.290
0.315
0.750
0.735
0.282
0.376
0.941
0.723
0.290
0.296
0.319
0.330
0.304
0.389
0.353
0.415
720
0.245
0.279 0.249
0.275
0.289
0.317
0.769
0.765
0.357
0.427
0.882
0.717
0.285
0.295
0.338
0.337
0.308
0.388
0.356
0.413
Avg
0.224
0.258
0.233 0.262
0.270
0.307
0.641
0.639
0.291
0.381
0.885
0.711
0.261
0.381
0.301
0.319
0.282
0.375
0.330
0.401
Electricity
96
0.146
0.238
0.148 0.240
0.181
0.270
0.219
0.314
0.193
0.308
0.201
0.317
0.169
0.273
0.168
0.272
0.247
0.345
0.197
0.282
192
0.160
0.252
0.162 0.253
0.188
0.274
0.231
0.322
0.201
0.315
0.222
0.334
0.182
0.286
0.184
0.289
0.257
0.355
0.196
0.285
336
0.176
0.270 0.178
0.269
0.204
0.293
0.246
0.337
0.214
0.329
0.231
0.338
0.200
0.304
0.198
0.300
0.269
0.369
0.209
0.301
720
0.221
0.306
0.225
0.317
0.246
0.324
0.280
0.363
0.246
0.355
0.254
0.361
0.222
0.321
0.220
0.320
0.299
0.390
0.245
0.333
Avg
0.176
0.267
0.178 0.270
0.205
0.290
0.244
0.334
0.214
0.327
0.227
0.338
0.193
0.296
0.192
0.295
0.268
0.365
0.212
0.300
Weather
96
0.168
0.208
0.174
0.214
0.177
0.218
0.158
0.230
0.217
0.296
0.266
0.336
0.173
0.223
0.172
0.220
0.221
0.306
0.196
0.255
192
0.216
0.253
0.221
0.254
0.225
0.259
0.206
0.277
0.276
0.336
0.307
0.367
0.245
0.285
0.219
0.261
0.261
0.340
0.237
0.296
336
0.276
0.296
0.278
0.296
0.278
0.297
0.272
0.335
0.339
0.380
0.359
0.395
0.321
0.338
0.280
0.306
0.309
0.378
0.283
0.335
720
0.352 0.348
0.358
0.347
0.354
0.348
0.398
0.418
0.403
0.428
0.419
0.428
0.414
0.410
0.365
0.359
0.377
0.427
0.345
0.381
Avg
0.253
0.276
0.258 0.278
0.259
0.281
0.259
0.315
0.309
0.360
0.338
0.382
0.288
0.314
0.259
0.287
0.292
0.363
0.265
0.317
Exchange
96
0.081
0.199
0.086
0.206
0.088
0.205
0.256
0.367
0.148
0.278
0.197
0.323
0.111
0.237
0.107
0.234
0.267
0.396
0.088
0.218
192
0.167
0.293
0.177
0.299 0.176 0.299
0.470
0.509
0.271
0.315
0.300
0.369
0.219
0.335
0.226
0.344
0.351
0.459
0.176
0.315
336
0.312 0.403
0.331
0.417
0.301
0.397
1.268
0.883
0.460
0.427
0.509
0.524
0.421
0.476
0.367
0.448
1.324
0.853
0.313
0.427
720
0.744
0.650
0.847
0.691
0.901
0.714
1.767
1.068
1.195
0.695
1.447
0.941
1.092
0.769
0.964
0.746
1.058
0.797
0.839
0.695
Avg
0.326
0.386
0.360
0.403
0.367
0.404
0.940
0.707
0.519
0.429
0.613
0.539
0.461
0.454
0.416
0.443
0.750
0.626
0.354
0.414
ETTm1
96
0.320
0.360
0.334
0.368
0.329 0.367
0.404
0.426
0.379
0.419
0.505
0.475
0.386
0.398
0.338
0.375
0.418
0.438
0.345
0.372
192
0.367
0.383
0.377
0.391
0.367
0.385
0.450
0.451
0.426
0.441
0.553
0.496
0.459
0.444
0.374
0.387
0.439
0.450
0.380
0.389
336
0.406 0.411
0.426
0.420
0.399
0.410
0.532
0.515
0.445
0.459
0.621
0.537
0.495
0.464
0.410
0.411
0.490
0.485
0.413
0.413
720
0459 0.447
0.491
0.459
0.454
0.439
0.666
0.589
0.543
0.490
0.671
0.561
0.585
0.516
0.478
0.450
0.595
0.550
0.474
0.453
Avg
0.388
0.400
0.407
0.410
0.387
0.400
0.513
0.496
0.448
0.452
0.588
0.517
0.481
0.456
0.400
0.406
0.485
0.481
0.403
0.407
ETTm2
96
0.176 0.260
0.180
0.264
0.175
0.259
0.287
0.366
0.203
0.287
0.255
0.339
0.192
0.274
0.187
0.267
0.286
0.377
0.193
0.292
192
0.245 0.307
0.250
0.309
0.241
0.302
0.414
0.492
0.269
0.328
0.281
0.340
0.280
0.339
0.249
0.309
0.399
0.445
0.284
0.362
336
0.305
0.343
0.311 0.348
0.305
0.343
0.597
0.542
0.325
0.366
0.339
0.372
0.334
0.361
0.321
0.351
0.637
0.591
0.369
0.427
720
0.404
0.399
0.412
0.407
0.402
0.400
1.730
1.042
0.421
0.415
0.433
0.432
0.417
0.413
0.408
0.403
0.960
0.735
0.554
0.522
Avg
0.283 0.327
0.288
0.332
0.281
0.326
0.757
0.610
0.305
0.349
0.327
0.371
0.306
0.347
0.291
0.333
0.571
0.537
0.350
0.401
ETTh1
96
0.383
0.398
0.386
0.405
0.414
0.419
0.423
0.448
0.376
0.419
0.449
0.459
0.513
0.491
0.384
0.402
0.654
0.599
0.386
0.400
192
0.433 0.435
0.441
0.436
0.460
0.445
0.471
0.474
0.420
0.448
0.500
0.482
0.534
0.504
0.436
0.429
0.719
0.631
0.437
0.432
336
0.464
0.452
0.487
0.458
0.501
0.466
0.570
0.546
0.459
0.465
0.521
0.496
0.588
0.535
0.491
0.469
0.778
0.659
0.481
0.459
720
0.485
0.484
0.503
0.491
0.500 0.488
0.653
0.621
0.506
0.507
0.514
0.512
0.643
0.616
0.521
0.500
0.836
0.699
0.519
0.516
Avg
0.441
0.442
0.454
0.447
0.469
0.454
0.529
0.522
0.440
0.460
0.496
0.487
0.570
0.537
0.458
0.450
0.747
0.647
0.456
0.452
ETTh2
96
0.306
0.351
0.297
0.349 0.302
0.348
0.745
0.584
0.358
0.397
0.346
0.388
0.476
0.458
0.340
0.374
0.707
0.621
0.333
0.387
192
0.371
0.394
0.380 0.400
0.388
0.400
0.877
0.656
0.429
0.439
0.456
0.452
0.512
0.493
0.402
0.414
0.860
0.689
0.477
0.476
336
0.415
0.430
0.428
0.432 0.426
0.433
1.043
0.731
0.496
0.487
0.482
0.486
0.552
0.551
0.452
0.452
1.000
0.744
0.594
0.541
720
0.425
0.442
0.427 0.445
0.431
0.446
1.104
0.763
0.463
0.474
0.515
0.511
0.562
0.560
0.462
0.468
1.249
0.838
0.831
0.657
Avg
0.379
0.404
0.383 0.407
0.387
0.407
0.942
0.684
0.437
0.449
0.450
0.459
0.526
0.516
0.414
0.427
0.954
0.723
0.559
0.515
$1^{st}$ Count
28
33
1
5
10 9
3
0
4
0
0
0
0
0
1
1
0
0
1
0